# How In-Context Learning Emerges from Training on Unstructured Data: On the Role of Co-Occurrence, Positional Information, and Noise Structures

### v1

@@ 2.2 In-context learning on dual-connected-relationship tasks
Let $K,L\geq 2$ and $S\geq 5$. Suppose we have infinitely many training sentences of length $S$, each comprising exactly two distinct $(c_{i},d_{i})$ pairs or $(c_{i},e_{i})$ pairs with uniform probability, and $S-4$ distinct $r_{i}$’s. We train a CBOW model with the squared loss and a large enough embedding dimension. Given a prompt $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$ ($c_{i_{1}}e_{i_{1}}\cdots c_{i_{\ell}}e_{i_{\ell}}c_{i_{\ell+1}}$) with distinct $i_{k}$’s, the model is more likely to predict $d_{i_{\ell+1}}$ ($e_{i_{\ell+1}}$) than $e_{i_{\ell+1}}$ ($d_{i_{\ell+1}}$).
Theorem 2 says that, when each training sentence includes two $(c_{i},d_{i})$ pairs or two $(c_{i},e_{i})$ pairs, a trained CBOW model is capable of task selection. To understand this result, consider the ICL prompt of the first type, i.e., $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$. Here, the output is more likely to be $d_{i_{\ell+1}}$ than $e_{i_{\ell+1}}$ since $d_{i_{\ell+1}}$ co-occurs with the other $d_{i_{j}}$’s in the training data (and $e_{i_{\ell+1}}$ does not). In Theorem 2, we require that each sentence contains either two distinct $(c_{i},d_{i})$ pairs or two distinct $(c_{i},e_{i})$ pairs for ICL to emerge. However, this may not be necessary as we empirically show next.
Experiments. We use the cross-entropy loss with $S=8$, $K=10$, $L=60$, and $\ell=3$. Each training sentence is equally likely to be a cd sentence (i.e., containing $(c_{i},d_{i})$ pairs) or a ce sentence (i.e., containing $(c_{i},e_{i})$ pairs), but not both. We explore multiple $(p_{0},p_{1},p_{2})$’s, where $p_{k}$ is the probability of having exactly $k$ pairs of $(c_{i},d_{i})$ for a cd sentence, or $k$ pairs of $(c_{i},e_{i})$ for a ce sentence.
@@
Additionally, we introduce three different scenarios: balanced, where all $L$ random words are equally likely to occur in both cd and ce sentences; imbalanced, where $L/3$ words are more likely to occur in cd (ce) sentences; and extreme, where $L/3$ of the words can only occur in cd (ce) sentences. Table 2 shows the accuracies of both tasks for each scenario, averaged over 10 repetitions. We observe a perfect accuracy when $(p_{0},p_{1},p_{2})\in\{(1/2,0/1,2),(0,1/2,1/2),(1/3,1/3,1/3)\}$ across all embedding dimensions and scenario types. The near-zero accuracy when $(p_{0},p_{1},p_{2})$ or $(0,1,0)$ or $(0,0,1)$ is again an artifact of the cross-entropy loss, as discussed in Section 2.1.
Interestingly, ICL works in the imbalanced and extreme scenarios when $(p_{0},p_{1},p_{2})=(1/2,1/2,0)$, where sentences do not contain more than one $(c_{i},d_{i})$ or $(c_{i},e_{i})$ pair. To see this, consider the balanced scenario where each $r_{i}$ is equally probable to appear in both types of sentences. Given a prompt of the form $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$, it is easy to see that the model should output $d_{i_{\ell+1}}$ or $e_{i_{\ell+1}}$ with equal probability. On the other hand, in the imbalanced and extreme scenarios, the signals from the $r_{i}$’s can allow for task selection, thus contributing to the success of ICL.
Table 3: ICL on dual-disconnected-relationship tasks, averaged over 10 repetitions, achieves perfect accuracy when $(p_{0},p_{1},p_{2})\in\{(1/2,0/1,2),(0,1/2,1/2),(1/3,1/3,1/3)\}$ regardless of architectures and embedding dimensions ($d_{E}$). When $(p_{0},p_{1},p_{2})=(1/2,1/2,0)$, ICL already performs well under the balanced scenario.
@@
2.3 In-context learning on dual-disconnected-relationship tasks
We next replicate the experiments in Section 2.2, but with two disconnected relationships $(c_{i},d_{i})$ and $(e_{i},f_{i})$. For example, $(c_{i},d_{i})$ might represent a country and its capital city and $(e_{i},f_{i})$ might represent a company and its CEO. The vocabulary consists of $c_{1:K},d_{1:K},e_{1:K},f_{1:K},r_{1:L}$, where $r_{i}$’s represent other words. Table 3 summarizes the accuracies of the ICL tasks $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$ and $e_{i_{1}}f_{i_{1}}\cdots e_{i_{\ell}}f_{i_{\ell}}e_{i_{\ell+1}}$ for each scenario, averaged over 10 repetitions. Similar to the connected setting in Section 2.2, we observe a perfect accuracy when $(p_{0},p_{1},p_{2})$ is one of $(1/2,0/1,2)$, $(0,1/2,1/2)$, or $(1/3,1/3,1/3)$ across all embedding dimensions and scenario types. However, when $(p_{0},p_{1},p_{2})=(1/2,1/2,0)$, ICL already works well in the balanced scenario. This is because the two relationships are disjoint, thus making task selection easier.
In addition, we consider a contaminated version of the training data where cd (ef) sentences can contain some $e_{i}$’s and $f_{i}$’s ($c_{i}$’s and $d_{i}$’s). We also obtain a perfect accuracy when $(p_{0},p_{1},p_{2})\in\{(1/2,0/1,2),(0,1/2,1/2),(1/3,1/3,1/3)\}$ across all embedding dimensions and scenario types.
2.4 Experiments on countries, US states, and their capital cities
We perform two experiments involving countries and their capital cities, as well as US states and their capital cities. The prompts follow the format $c_{1}d_{1},c_{2}d_{2},\cdots,c_{6}d_{6},c_{7}$, where $c_{i}$ is a country or US state and $d_{i}$ is its capital city. Using the LLaMA 2 model [Touvron et al., 2023], we compare the prediction for each prompt with its corresponding $d_{7}$. The experimental results support the theory.
In the first experiment, we focus on 160 countries with a population exceeding one million in 2022. Among these countries, 31 have capital cities that are not their most populous cities, denoted by type A. The remaining 129 countries fall under type B. Each ICL prompt includes three type A countries among $c_{1},\cdots,c_{6}$ to emphasize that the desired relationship is (country)-(capital) rather than (country)-(largest city). Subsequently, we randomly generate 1,000 prompts, with 500 having a $c_{7}$ representing a type A country and 500 having a $c_{7}$ representing a type B country. The ICL accuracies corresponding to type A and type B prompts are $0.58$ and $0.96$, respectively.
In the second experiment, we consider all 50 states, among which 33 are of type A and 17 are of type B, defined similarly. The ICL accuracies corresponding to type A and type B prompts are found to be $0.69$ and $0.84$, respectively. From both experiments, we notice that LLaMA 2 performs better on type B prompts (i.e., the capital city as the largest city). This suggests that ICL may arise from co-occurrence information, as larger cities tend to appear more frequently compared to smaller ones.
@@ 2.5 Experiments on a synthetic corpus
We conduct experiments on a synthetic corpus consisting of (country)-(capital) and (country)-(IOC code) relationships. Each sentence in the corpus is categorized into exactly one of six possible categories: (1) exactly one country-capital pair; (2) exactly two country-capital pairs; (3) exactly one country-IOC pair; (4) exactly two country-IOC pairs; (5) exactly one country without any pair; and (6) no country. In sentences with country-capital pairs, each capital city can appear in any position relative to the country. Conversely, in sentences with country-IOC pairs, each IOC code must directly follow the country. The corpus generation process is as follows:
  1. 1.
  2. 2.
  3. 3.
    Example: Gabon (GAB) protects its diverse rainforests and wildlife.
  4. 4.
  5. 5.
    Generate 60 sentences without any country, capital city, or IOC code.
  6. 6.
  7. 7.
Two models are trained on this corpus: a CBOW and a five-layer two-head autoregressive transformer. Both models have an embedding dimension of $100$. We then compare the ICL accuracies for both relationships given one to five in-context examples. For the CBOW model, the country-capital accuracies are $(0.77,0.66,0.67,0.65,0.63)$ and the country-IOC accuracies are $(0.08,0.17,0.16,0.23,0.34)$. Here, the $i$-th number corresponds to the accuracy given $i$ in-context examples. For the transformer, the accuracies are $(0.05,0.26,0.04,0.02,0.09)$ and $(0.57,0.38,0.91,0.85,0.83)$, respectively.
@@
In this section, we examine another common example of in-context learning (ICL), where the task involves predicting the first (or second) token given a sequence of tokens. To understand the significance of positional information (unlike the tasks in Section 2), we consider a simpler task: modeling sequences of tokens in the form $x_{i_{1}}x_{i_{2}}x_{i_{3}}x_{i_{1}}$. Theorem 3 underscores the necessity of incorporating positional information to correctly predict $x_{i_{1}}$ from $x_{i_{1}}x_{i_{2}}x_{i_{3}}$ in a single-layer model, and provides a construction of a basic attention-based model capable of achieving zero loss and perfect accuracy on this task. Its proof is in Appendix C.
@@
Here, $f(\{x_{i_{1}},x_{i_{2}}\},x_{i_{3}})$ represents a scenario where the model lacks positional information (e.g., $f$ is a one-layer autoregressive transformer without positional embeddings). In this scenario, the output of this function is identical for inputs $x_{i_{1}}x_{i_{2}}x_{i_{3}}$ and $x_{i_{2}}x_{i_{1}}x_{i_{3}}$, which leads to the impossibility of attaining zero loss. In contrast, $\tilde{f}(\{(x_{i_{1}},1),(x_{i_{2}},2)\},(x_{i_{3}},3))$ refers to a scenario where the model has access to positional information. We provide a construction of $\tilde{f}$ that achieves zero loss in Appendix C.
Experiments. We validate Theorem 3 by training transformers to autoregressively learn sequences of the form $x_{i_{1}}x_{i_{2}}x_{i_{3}}x_{i_{1}}$, and assessing their accuracy in predicting the last token on a separate test data of the same pattern. We use $|V|=20$ and an embedding dimension of $10$. We consider these settings: (i) number of layers: 1, 5; (ii) positional embeddings: learned, sinusoidal, no positional embeddings; and (iii) train-test split: each token in the vocabulary is the first token in both the training and test sets (Both), each token in the vocabulary is the first token in either set, but not both (Either).
Table 4 summarizes the results. Two main findings emerge: (1) for the model to successfully generalize to unseen sentences, each token in $\mathcal{V}$ should be present as the first token in both the training and test sets; (2) positional embeddings are crucial when using only one attention layer.
Multiple layers. With multi-layer models, positional information can be encoded without explicit positional embeddings. This is summarized in Proposition 4, whose proof is in Appendix D.
Proposition 4 (Multi-layer models can encode positions).
Proposition 4 shows that we generally have $t(x_{i_{1}}x_{i_{2}}x_{i_{3}})\neq t(x_{i_{2}}x_{i_{1}}x_{i_{3}})$, unlike in the one-layer case. Consequently, high accuracy is achievable without positional embeddings, as shown in Table 4.
Table 4: Prediction accuracy with single/multi-layer models. For successful ICL, it is crucial that the first token of sentences in the training set covers the entire vocabulary (Both). In this case, positional embeddings are essential, especially when using a one-layer model.
@@
Roadmap of Section 3. In the rest of this section, we consider settings where each sentence contains repeating patterns. Section 3.1 focuses on a simple scenario where training sentences follow the form abacdc, where $a\neq b$ and $c\neq d$, or a noisy variation of it. The ICL prompts maintain the same pattern but use different combinations of ab and cd from those in the training data. The goal is to understand what types of training data facilitate ICL in clean or noisy scenarios. Section 3.2 explores a more realistic case where two possible patterns are present: repeating the first letter (abca) and repeating the second letter (abcb).
3.1 In-context learning on single-pattern tasks
In this section, we examine the case where the training sentences follow a specific pattern of the form abacdc. Also, we analyze the impact of adding noise tokens to the training sentences on the ICL ability of autoregressive models. To formalize the discussion, let the vocabulary be $\mathcal{V}\cup\mathcal{N}$, where $\mathcal{N}$ represents the noise tokens. Define $S=\{(a,b)\leavevmode\nobreak\ |\leavevmode\nobreak\ a,b\in\mathcal{V},a\neq b\}$ and partition $S$ into $S_{1}$ and $S_{2}$, where $\{c[1]\leavevmode\nobreak\ |\leavevmode\nobreak\ c\in S_{1}\}=\{c[1]\leavevmode\nobreak\ |\leavevmode\nobreak\ c\in S_{2}\}=\mathcal{V}$ and $c[i]$ denotes the $i$-th element of $c$, to ensure training sentences are distinct from the ICL prompts. Consider three different scenarios:
  1. 1.
  2. 2.
    One-noisy: Training data follow the form abacdc where $ab,cd\in S_{1}$, with one noise token $n\in\mathcal{N}$ randomly inserted anywhere except the last position (to ensure ICL prompts do not resemble the training data). ICL prompts follow the form abacd where ab, cd $\in S_{2}$.
  3. 3.
    Block-noisy: Training data follow the form abacdc where $ab,cd\in S_{1}$, with three consecutive noise tokens $n_{1},n_{2},n_{3}\in\mathcal{N}$ randomly inserted while preserving the aba and cdc blocks. ICL prompts follow the form abacdcef where ab, cd, ef $\in S_{2}$.
We set the vocabulary size $|V|=20$, the number of noise tokens $N=20$, and use only one attention layer as additional layers do not improve performance. Table 5 reveals interesting phenomena. Firstly, under the clean data scenario, ICL performs exceptionally well, with an observed performance increase with learned positional embeddings and a larger embedding dimension. However, ICL is notably challenging under the one-noisy scenario. In the block-noisy scenario, learned positional embeddings are crucial for satisfactory ICL performance. Theorem 5 formalizes these findings.
Theorem 5 (Blocked noise structure facilitates ICL).
The proof is in Appendix E. Theorem 5 says that ICL works perfectly under the block-noisy scenario, yet fails to work under the one-noisy scenario. However, as shown in Table 5, the use of sinusoidal positional embeddings significantly enhances prediction accuracy in the one-noisy scenario. This may be due to the fact that sinusoidal embeddings can encode relative positional information [Vaswani et al., 2017]. For example, training sentences of the form nabacdc, where n $\in\mathcal{N}$, may help in predicting the most likely token following the ICL prompt abacd.

### v2

@@ 2.2 In-context learning on dual-connected-relationship word analogy tasks
Let $K,L\geq 2$ and $S\geq 5$. Suppose each training sentence of length $S$ is generated by selecting two distinct $(c_{i},d_{i})$ pairs or $(c_{i},e_{i})$ pairs, and $S-4$ distinct $r_{i}$’s uniformly at random. We train a CBOW model with the squared loss and a large enough embedding dimension. Given a prompt $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$ ($c_{i_{1}}e_{i_{1}}\cdots c_{i_{\ell}}e_{i_{\ell}}c_{i_{\ell+1}}$) with distinct $i_{k}$’s, the model is more likely to predict $d_{i_{\ell+1}}$ ($e_{i_{\ell+1}}$) than $e_{i_{\ell+1}}$ ($d_{i_{\ell+1}}$). (The proof is in Appendix D.)
According to Theorem 2, when each training sentence includes two $(c_{i},d_{i})$ pairs or two $(c_{i},e_{i})$ pairs, a trained CBOW model is capable of performing task selection. To intuitively understand this result, consider the ICL prompt of the first type, i.e., $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$. Here, the output is more likely to be $d_{i_{\ell+1}}$ than $e_{i_{\ell+1}}$ since $d_{i_{\ell+1}}$ co-occurs with the other $d_{i_{j}}$’s in the training data (and $e_{i_{\ell+1}}$ does not). Note that in Theorem 2, we unrealistically require each sentence to contain either two distinct $(c_{i},d_{i})$ pairs or $(c_{i},e_{i})$ pairs. However, this condition is not necessary as we empirically show next.
Experiments. We use the cross-entropy loss with $S=8$, $K=10$, $L=60$, and $\ell=3$. Each training sentence is equally likely to be a cd sentence (i.e., containing $(c_{i},d_{i})$ pairs) or a ce sentence (i.e., containing $(c_{i},e_{i})$ pairs), but not both. We explore multiple $(p_{0},p_{1},p_{2})$’s, where $p_{k}$ is the probability of having exactly $k$ pairs of $(c_{i},d_{i})$ for a cd sentence, or $k$ pairs of $(c_{i},e_{i})$ for a ce sentence. Additionally, we introduce three different scenarios: balanced, where all $L$ random words are equally likely to occur in both cd and ce sentences; imbalanced, where $L/3$ words are more likely to occur in cd (ce) sentences; and extreme, where $L/3$ of the words can only occur in cd (ce) sentences. More details are provided in Appendix J.
@@
Results. Table 2 shows the accuracies of both tasks for each scenario, averaged over 10 repetitions. We observe a perfect accuracy when $(p_{0},p_{1},p_{2})\in\{(1/2,0/1,2),(0,1/2,1/2),(1/3,1/3,1/3)\}$ across all embedding dimensions and scenario types. The near-zero accuracy when $(p_{0},p_{1},p_{2})$ or $(0,1,0)$ or $(0,0,1)$ is again an artifact of the cross-entropy loss, as discussed in Section 2.1.
Interestingly, ICL works in the imbalanced and extreme scenarios when $(p_{0},p_{1},p_{2})=(1/2,1/2,0)$, where sentences do not contain more than one $(c_{i},d_{i})$ or $(c_{i},e_{i})$ pair. To see this, consider the balanced scenario where each $r_{i}$ is equally probable to appear in both types of sentences. Given a prompt of the form $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$, it is easy to see that the model should output $d_{i_{\ell+1}}$ or $e_{i_{\ell+1}}$ with equal probability. On the other hand, in the imbalanced and extreme scenarios, the information from the $r_{i}$’s can allow for task selection, thus contributing to the success of ICL.
2.3 In-context learning on dual-disjoint-relationship tasks
We next replicate the experiments in Section 2.2, but with disjoint word pair relationships of two distinct types with no overlapping tokens, i.e., $(c_{i},d_{i})$ and $(e_{i},f_{i})$. For example, $(c_{i},d_{i})$ represents a country and its capital city, and $(e_{i},f_{i})$ represents a company and its CEO. Our vocabulary consists of $c_{1:K},d_{1:K},e_{1:K},f_{1:K},r_{1:L}$, where $r_{i}$’s represent other words; see Appendix J for details.
Results. Table 3 presents the accuracies of the ICL tasks $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$ and $e_{i_{1}}f_{i_{1}}\cdots e_{i_{\ell}}f_{i_{\ell}}e_{i_{\ell+1}}$ for each scenario, averaged over 10 repetitions. Similar to the connected setting in Section 2.2, we observe a perfect accuracy when $(p_{0},p_{1},p_{2})\in\{(1/2,0/1,2),(0,1/2,1/2),(1/3,1/3,1/3)\}$ across all embedding dimensions and scenario types. However, when $(p_{0},p_{1},p_{2})=(1/2,1/2,0)$, ICL already works well in the balanced scenario. Intuitively, this is because the two relationships are disjoint, thus making task selection easier.
In addition, we consider a contaminated version of the training data where cd (ef) sentences can contain some $e_{i}$’s and $f_{i}$’s ($c_{i}$’s and $d_{i}$’s). We also obtain a perfect accuracy when $(p_{0},p_{1},p_{2})$ is in $\{(1/2,0/1,2),(0,1/2,1/2),(1/3,1/3,1/3)\}$ across all embedding dimensions and scenario types.
Table 3: ICL on dual-disjoint-relationship tasks, averaged over 10 repetitions, achieves perfect accuracy when $(p_{0},p_{1},p_{2})\in\{(1/2,0/1,2),(0,1/2,1/2),(1/3,1/3,1/3)\}$ regardless of architectures and embedding dimensions ($d_{E}$). When $(p_{0},p_{1},p_{2})=(1/2,1/2,0)$, ICL already performs well under the balanced scenario.
@@ 2.4 Experiments on a synthetic corpus
We conduct experiments on a synthetic corpus consisting of (country)-(capital) and (country)-(IOC code) relationships. Each sentence in the corpus is categorized into exactly one of six possible categories: (1) exactly one country-capital pair; (2) exactly two country-capital pairs; (3) exactly one country-IOC pair; (4) exactly two country-IOC pairs; (5) exactly one country without any pair; and (6) no country. In sentences with country-capital pairs, each capital city can appear in any position relative to the country. Conversely, in sentences with country-IOC pairs, each IOC code must directly follow the country. The data source and corpus generation process are detailed in Appendix J.
Two models are trained on this corpus: a CBOW and a five-layer two-head autoregressive transformer. Both models have an embedding dimension of $100$. We then compare the ICL accuracies for both relationships given one to five in-context examples. For the CBOW model, the country-capital accuracies are $(0.81,0.82,0.78,0.73,0.65)$ and the country-IOC accuracies are $(0.15,0.38,0.59,0.71,0.79)$. Here, the $i$-th number corresponds to the accuracy given $i$ in-context examples. For the transformer, the accuracies are $(0.00,0.15,0.34,0.22,0.07)$ and $(1.00,0.77,0.78,0.97,0.99)$, respectively.
@@
We examine another common example of in-context learning (ICL), where the task involves predicting the first (or second) token in a sequence. This task resembles general logic reasoning tasks that require recognizing patterns that do not typically co-occur in a sentence, such as (word)-(first letter) [62, 13]. While Section 2 shows that positional encoding is irrelevant for ICL in word analogy tasks, positional information proves essential for such logic reasoning tasks. Specifically, we consider a simpler task of modeling $x_{i_{1}}x_{i_{2}}x_{i_{3}}x_{i_{1}}$. Theorem 3 underscores the importance of positional information to correctly predict $x_{i_{1}}$ from $x_{i_{1}}x_{i_{2}}x_{i_{3}}$ in a single-layer model, and provides a construction of an attention-based model achieving zero loss and perfect accuracy on this task. Its proof is in Appendix E.
@@
Here, $f(\{x_{i_{1}},x_{i_{2}}\},x_{i_{3}})$ represents a scenario where the model lacks positional information (e.g., $f$ is a one-layer autoregressive transformer without positional embeddings). Note that the output of this function is identical for inputs $x_{i_{1}}x_{i_{2}}x_{i_{3}}$ and $x_{i_{2}}x_{i_{1}}x_{i_{3}}$, which leads to the impossibility of attaining zero loss. In contrast, $\tilde{f}(\{(x_{i_{1}},1),(x_{i_{2}},2)\},(x_{i_{3}},3))$ refers to a scenario where the model has access to positional information. We provide a construction of $\tilde{f}$ that achieves zero loss in Appendix E.
Experiments. We validate Theorem 3 by training transformers with causal masking to autoregressively learn sequences of the form $x_{i_{1}}x_{i_{2}}x_{i_{3}}x_{i_{1}}$, and assessing their accuracy in predicting the last token on a separate test data of the same pattern. We use $|V|=20$ and an embedding dimension of $10$. We consider these settings: (i) number of layers: 1, 5; (ii) positional embeddings: learned, sinusoidal, no positional embeddings; and (iii) train-test split: each token in the vocabulary is the first token in both the training and test sets (Both), each token in the vocabulary is the first token in either set, but not both (Either). More details are provided in Appendix J.
Table 4: Prediction accuracy with single/multi-layer models. For ICL to occur, the first tokens of training sentences should cover the entire vocabulary (Both). Also, positional embeddings are essential, especially in one-layer models.
@@
Results. Table 4 summarizes the results. Two main findings emerge: (1) for the model to generalize to unseen sentences, each token in $\mathcal{V}$ should be present as the first token in both the training and test sets; (2) positional embeddings are crucial when using only one attention layer. Note that in practice, the condition in (1) is likely met due to the vast size of LLMs’ pre-training data.
Multiple layers. Proposition 4 shows that multi-layer models can encode positional information without explicit positional embeddings.
Proposition 4 (Multi-layer models can encode positions).
The proof is in Appendix F. Proposition 4 shows that we generally have $t(x_{i_{1}}x_{i_{2}}x_{i_{3}})\neq t(x_{i_{2}}x_{i_{1}}x_{i_{3}})$, unlike in the one-layer case. Consequently, high accuracy is achievable without positional embeddings, as shown in Table 4. This result parallels findings in Haviv et al. [26] that autoregressive transformers implicitly encode positions, even without positional embeddings.
Roadmap of Section 3. In the rest of this section, we consider settings where each sentence contains repeating patterns. Section 3.1 focuses on a simple scenario where training sentences follow the form abacdc, where $a\neq b$ and $c\neq d$, or a noisy variation of it. The ICL prompts maintain the same pattern but use different combinations of ab and cd from those in the training data. Our goal is to understand what types of training data facilitate ICL in clean or noisy scenarios. Section 3.2 explores a more realistic case where two possible patterns are present: repeating the first letter (abca) and repeating the second letter (abcb).
@@
RoPE [48] 0.31 0.00 0.03 0.48 0.00 0.00
3.1 In-context learning on single-pattern tasks
In this section, we examine the case where the training sentences follow the pattern abacdc. To replicate real-world training scenarios, we also analyze how incorporating nuisance tokens into the training sentences affects the ICL capability of autoregressive models. To formalize the discussion, let our vocabulary be $\mathcal{V}\cup\mathcal{N}$, where $\mathcal{N}$ represents the nuisance tokens. We define $S=\{(a,b)\leavevmode\nobreak\ |\leavevmode\nobreak\ a,b\in\mathcal{V},a\neq b\}$ and partition $S$ into $S_{1}$ (for training sentences) and $S_{2}$ (for ICL prompts). This is to ensure that training sentences are distinct from ICL prompts. Furthermore, we assume $\{c[1]\leavevmode\nobreak\ |\leavevmode\nobreak\ c\in S_{1}\}=\{c[1]\leavevmode\nobreak\ |\leavevmode\nobreak\ c\in S_{2}\}=\mathcal{V}$, where $c[i]$ is the $i$-th element of $c$. In other words, each token in $\mathcal{V}$ can be the first token in both the training sentences and ICL prompts. We consider three scenarios:
  1. 1.
  2. 2.
    One-noisy: Training data follow the form abacdc where $ab,cd\in S_{1}$, with one nuisance token $n\in\mathcal{N}$ randomly inserted anywhere except the last position (to ensure ICL prompts do not resemble the training data). ICL prompts follow the form abacd where ab, cd $\in S_{2}$.
  3. 3.
    Block-noisy: Training data follow the form abacdc where $ab,cd\in S_{1}$, with three consecutive nuisance tokens $n_{1},n_{2},n_{3}\in\mathcal{N}$ randomly inserted while preserving the aba and cdc blocks. ICL prompts follow the form abacdcef where ab, cd, ef $\in S_{2}$.
We set the vocabulary size $|V|=20$, the number of nuisance tokens $N=20$, and use only one attention layer as additional layers do not improve performance. See Appendix J for more details.
Results. Table 5 reveals interesting phenomena. First, under the clean data scenario, ICL performs exceptionally well, with an observed performance increase with learned positional embeddings and a larger embedding dimension. However, ICL is notably challenging under the one-noisy scenario. In the block-noisy scenario, learned positional embeddings are crucial for satisfactory ICL performance. Theorem 5 formalizes these findings.
Theorem 5 (Blocked nuisance token structure facilitates ICL).
The proof is in Appendix G. Theorem 5 says that ICL works perfectly under the block-noisy scenario, yet fails to work under the one-noisy scenario. However, as shown in Table 5, the use of sinusoidal positional embeddings significantly enhances prediction accuracy in the one-noisy scenario. This may be due to the fact that sinusoidal embeddings can encode relative positional information [53]. For example, training sentences of the form nabacdc, where n $\in\mathcal{N}$, may help in predicting the most likely token following the ICL prompt abacd.